Sol Summary – February 12th

Crew 292 Sol Summary Report 12-02-2024

Sol: Sol #8

Summary Title: Business on Mars as usual

Author’s name: Annalea Beattie

Mission Status: Nominal

Sol Activity Summary: Even after great excitement, a late night and a terrific outreach interview run by Sakshi Sharma in India, we completed two successful EVA’s today to Kissing Camel Ridge using Spirit and Perseverance rovers. In our two groups we covered 800 metres of Kissing Camel Ridge, up and down. There is still ice on the Morrison even though it was a beautiful day. I kept an eye out for terrestrial mountain lions and saw nothing even though paw prints have been spotted near the road. EVA’s #9 and # 10 focused on inverted river channels and different types of weathering including cross-bedded sandstone. These complex signs of ancient fluvio-lacustrine dunes are very similar to large scale outcrops photographed by Curiosity on Mars (and they are amazing to draw). Both EVA’s today ran like clockwork and we continue to achieve our science goals.

The last couple of days especially, our joint Australian Indian crew is in its full stride. Morale is great. This crew is a considerate crew and we are enjoying being a team. Every day the group at home cares for the group away, ensuring jobs are done and food is on the table when the others return and this is what happened today. In the hab there is a rhythm to our day as if this is the way we have always lived. We have really learnt in the last week to pull together to conserve our resources. This includes, for instance, portioning food so there is no waste. This is a challenge when cooking a meal for seven hearty vegetarians but we are getting the hang of it. Today during EVA# 9 the home group cleaned and cooked and discussed our crew plans for analogue research and astrobiology in India. We decided that it’s time we reinvigorated Mars Society India and our geologist Bharti Sharma has taken the lead. In the afternoon while our second crew were out on EVA #10, Clare assisted Daniel with DNA extractions in the Science Dome, Aditya worked on his image of a spiral galaxy and Mehnaz worked on her experiments in the Green Hab checking plants for rates of evapotranspiration. Chocolate orange peanut butter brownies were cooked by Dan and a sabji dinner with chapattis is now ready for us after Comms.

Look Ahead Plan: One EVA requested tomorrow to Candor Chasma and Compass Rock

Anomalies in work: None

Weather: Cold but sunny all day with clear skies.

Crew Physical Status: Nominal

EVA: All crew in two groups (EVA #9 and EVA #10) to the east and then west of Kissing Camel Ridge.

Reports to be filed: Sol Summary, Journalist Report, Greenhab Report, Operations Report, EVA report, Astronomy Report, EVA request.

Support Requested: None

Copyright © The Mars Society. All rights reserved. | Main Site